# Many-Body Localization in the Age of Classical Computing

### v2

@@ IV.4 Conclusion
In the following section V, we briefly outline numerical approaches typically employed in studies of ergodicity and MBL. A reader interested in physics of disordered quantum many-body may skip Sec. V, and proceed directly to Sec. VI, in which we review numerical investigations of the ETH-MBL crossover in disordered many-body systems, providing concrete examples of the quantities which follow the general patterns described in the present section.
@@ V.1 Exact diagonalization
Exact diagonalization (ED) is a numerical method in which a certain number of eigenvalues $E_{m}$ and eigenvectors |$\ket{m}$ of Hamiltonian $\hat{H}$ of a many-body system is calculated. To that end, one needs to specify the matrix $H$ corresponding to the Hamiltonian $\hat{H}$ in a selected many-body basis, as described, e.g., in an introductory fashion in [732, 303]. Once the Hamiltonian matrix $H$ is calculated, one may employ a full exact diagonalization routine, using one of the available linear algebra packages, for instance, the LAPACK package [26], to compute all eigenvalues $E_{m}$ and eigenvectors |$\ket{m}$ of $\hat{H}$. Subsequently, the questions about thermalization (1) in the many-body system described by the Hamiltonian $\hat{H}$ can be answered. However, understanding whether the investigated system is in the MBL phase, (6), or whether it follows the ETH requires taking the large system size limit $L$. The computational resources required by the full ED quickly become prohibitive with increasing $L$ due to the exponential growth of the Hilbert space dimension $\mathcal{N}$ with the system size $L$, impeding the extrapolation of the results to $L\to\infty$. The typical matrix size accessible for full ED calculations in present-day computational studies is of the order of $\mathcal{N}\sim 10^{5}$. Hence, full ED calculations for the disordered XXZ spin chain (7) in the $\sum_{i}S^{z}_{i}=0$ sector are restricted to $L\leq 18$. Note that investigations of MBL usually involve disorder average, which makes studies of larger system sizes infeasible.
@@
An observation that enables the exploration of system sizes beyond the reach of full ED is that the Hamiltonians of quantum many-body systems can often be written as sparse matrices for an appropriately chosen basis. The number of non-zero off-diagonal entries per row in such sparse matrices is much smaller than the matrix size $\mathcal{N}$. For example, the Hamiltonian of the disordered XXZ spin chain (7), written in the eigenbasis of $S^{z}_{i}$ operators, contains no more than $L$ non-zero off-diagonal elements per row. By retaining the non-vanishing elements of the sparse Hamiltonian matrix $H$ solely, it becomes possible to significantly reduce the computation time associated with matrix-vector multiplication in comparison to a dense matrix (which, by definition, has ${}\mathcal{N}$ non-zero entries per row). This opens up the possibility of using Lanczos algorithm [353], which employs repeated matrix-vector multiplications to compute eigenvectors corresponding to eigenvalues at the edges of the spectrum of $H$. The Lanczos algorithm’s convergence slows considerably as the number of requested eigenvectors at the spectrum edges increases. The Lanczos algorithm needs to be paired with a so-called spectral transformation to make it practical to calculate the eigenvectors in the middle of the spectrum of $H$ that are of interest in studies of MBL.
Shift-and-invert method relies on a transformation of the Hamiltonian matrix $H$ via
$H\rightarrow R_{\sigma}(H)=(\sigma-H)^{-1},$ (21)
@@
where $\sigma$ is a chosen energy target. The spectral transformation (21) has two important properties: i) the eigenvectors |$\ket{m}$ of $H$ are still eigenvectors of the transformed matrix $R_{\sigma}(H)$, ii) the eigenvalues are mapped according to $E_{m}\rightarrow R_{\sigma}(E_{m})$, which, by the construction of the function, ensures that the eigenvalues in the vicinity of the energy target $\sigma$ become the eigenvalues at the spectrum edges of the transformed matrix $R_{\sigma}(H)$. The two properties imply that the Lanczos algorithm for the transformed matrix $R_{\sigma}(H)$ converges to eigenvalues of $H$ close to $\sigma$. Shift-and-invert proved to be fruitful in studies of MBL [393], extending to $L\leq 26$ the range of system sizes for which calculation of eigenvectors in the middle of the spectrum of (7) is possible. The main difficulty of the shift-and-invert method is the presence of the inverse operator in (21). For interacting many-body systems, the Lanczos iteration with $R_{\sigma}(H)$ can be performed efficiently by a massively parallel $LU$ decomposition [24, 271, 42]. However, the sparsity pattern of the Hamiltonian matrix $H$ is not preserved during the $LU$ decomposition, which leads to a huge memory overhead, and was identified as the main bottle-neck of the shift-and-invert method [484].
@@
Polynomial spectral transformations [541], which replace the complicated function (21) by a polynomial of $H$, constitute an alternative to the shift-and-invert method. Calculation of eigenvectors in the middle of the spectrum of a large sparse matrix such as (7) can be efficiently carried out with an algorithm POLFED [597] that relies on the following polynomial spectral transformation
$H\rightarrow P^{K}_{\sigma}(H)=\frac{1}{D}\sum_{k=0}^{K}c^{\sigma}_{k}T_{k}(H),$ (22)
@@
where $T_{k}(x)$ denotes $k$-th Chebyshev polynomial, the coefficients read $c^{\sigma}_{k}=\sqrt{4-3\delta_{0,k}}\cos(n\arccos{\sigma})$, and $D$ is a normalization factor. For sufficiently high and appropriately chosen order $K$, the polynomial $P^{K}_{\sigma}(E)$ has a sharp peak around the energy target $\sigma$ since the coefficients $c^{\sigma}_{n}$ are obtained from expanding a Dirac delta function centered at this $\sigma$ value. Hence, the Lanczos algorithm for $P^{K}_{\sigma}(H)$ converges to eigenvectors of $H$ with energies around the chosen energy target. Notably, besides basic linear algebra operations, the Lanczos algorithm employs only the multiplication of vectors by $P^{K}_{\sigma}(H)$. The multiplication of vectors by $P^{K}_{\sigma}(H)$ reduces to repeated matrix-vector multiplications by the matrix $H$ and linear algebra operations. Hence, the sparse structure of $H$ is preserved by the POLFED algorithm, which significantly reduces the procedure’s memory cost. For instance, POLFED requires $\approx 100$ GB of RAM to calculate $2000$ eigenvectors in the middle of the spectrum of (7) for $L=24$, easily fitting into a single computing node of a present-day supercomputer. The shift-and-invert method requires $\approx 10^{4}$ GB of RAM for this task. At the same time, the total CPU time used by the two methods is of similar order; see [597] and [484] for further benchmarks. In passing, we note that combining a spectral transformation based on a second-order polynomial of $H$ with appropriate preconditioning methods provides another exact diagonalization approach to MBL [53].
@@
Finally, the polynomial spectral transformations can also be used for efficient calculations of eigenvectors for large Floquet operator $U_{F}$ [385]. In that case, analogously to the POLFED algorithm, one considers a high order polynomial of the Floquet operator,
$g_{K}(U_{F})=\sum_{m=0}^{K}e^{-im\phi_{\mathrm{tg}}}U_{F}^{m},$ (23)
@@
where $\phi_{\mathrm{tg}}$ is the target eigenphase and the order of polynomial $K$ is fixed as $K=f\frac{\mathcal{N}}{N_{\mathrm{ev}}}$, where $N_{\mathrm{ev}}$ is the number of requested eigenvectors, the Hilbert space dimension is $\mathcal{N}=2^{L}$, and the factor $f=1.46$ was obtained from an optimization of the performance of the algorithm. The Arnoldi iteration [33] for the operator $g_{K}(U_{F})$, yields $N_{\mathrm{ev}}$ eigenvectors |$\ket{\psi_{n}}$ corresponding to the dominant eigenvalues of $g_{K}(U_{F})$ (which is no longer a unitary operator). The latter, by construction, correspond to eigenvectors of $U_{F}$ to eigenphases closest to $\phi_{\mathrm{tg}}$. The convergence is reached after approximately $\alpha N_{\mathrm{ev}}$ steps of the iteration, where $\alpha\approx 2.1$. This algorithm is especially effective when the matrix-vector multiplication by the operator $U_{F}$ can be efficiently implemented. This is the case for the Kicked Ising model, as described in Sec. X.4, or for various types of quantum circuits, one example was considered by [427]. For these cases, on present day computers, the algorithm enables study of eigenphases and eigenvectors of Floquet operators for $L=20$ or more (corresponding to matrix size $\mathcal{N}>10^{6}$), while the full exact diagonalization of Floquet operators is feasible up to $L\approx 15$ and $\mathcal{N}\approx 3\cdot 10^{4}$. In general, efficient matrix vector products can be formulated with a matrix product operator formulation of $U_{F}$ [731].
V.2 Time evolution
@@
Exact diagonalization approaches allow us to study eigenvectors of $\hat{H}$, which encode, via (2), the limit of infinite times, $t\to\infty$. Moreover, the complete information about the eigenvalues and eigenvectors enables calculation of the time evolved state |$\ket{\psi(t)}$, and consequently, the time evolution of any observable $\hat{A}$ with (1). The memory complexity of the full ED is $O(\mathcal{N}^{2})$, while the space needed to store the time evolved state |$\ket{\psi(t)}$ scales linearly with $\mathcal{N}$. This observation, together with the sparsity of the Hamiltonian matrix $H$, motivates the search for more effective time evolution algorithms.
@@
The efficient algorithm of time propagation of Schrödinger equation was proposed in [637]. The algorithm is based on expansion of the time evolution operator $U(\Delta t)=e^{-iH\Delta t}$ over a time step $\Delta t$ into Chebyshev polynomials $T_{k}(x)$ of the Hamiltonian matrix
$U(\Delta t)\approx\mathrm{e}^{-\mathrm{i}b\Delta t}\left(J_{0}(a\Delta t)+2\sum_{k=1}^{N}(-i)^{k}J_{k}(a\Delta t)T_{k}\left(\mathcal{H}\right)\right),$ (24)
@@
where $a=(E_{\rm max}-E_{\rm min})/2$, $b=(E_{\rm max}+E_{\rm min})/2$ and $E_{\rm min}$ ($E_{\rm max}$) is the energy of the ground state (the highest excited eigenstate) of $H$. The Hamiltonian is rescaled as $\mathcal{H}=\frac{1}{a}(H-b)$, and $J_{k}(t)$ is the Bessel function of the order $k$. The order of expansion, $N\ll\mathcal{N}$, needed to assure convergence of (24) up to any desired precision, can be determined numerically for a given time step $\Delta t$, see, for instance, [604]. An application of $U(\Delta t)$ allows to update the state of the system |$\ket{\psi{(t_{0}+\Delta t)}}=U(\Delta t)\ket{\psi{(t_{0})}}$, and, upon repetition, results in the time-evolved state |$\ket{\psi(t)}$ with $t>t_{0}$. Since, for a fixed order $N$, $U(\Delta t)$ is simply a polynomial of the matrix $H$, the calculation of $U(\Delta t)\ket{\psi{(t_{0})}}$ reduces to multiple multiplications of vectors by matrix $H$ and linear algebra operators. Hence, the efficiency of the method relies crucially on the sparsity of the matrix $H$. This Chebyshev time evolution algorithm allows to calculate the time-evolved state |$\ket{\psi(t)}$ couple of times faster than Suzuki-Trotter-based schemes [170], see also [202].
@@
A closely related method, called Krylov time evolution, was proposed in [444]. A direct truncation of the Taylor expansion of the time evolution operator $U(\Delta t)\ket{\psi{(t_{0})}}=\sum_{k}\frac{(-i\Delta t)^{k}}{k!}H^{k}\ket{\psi{(t_{0})}}$ to order $m$ is numerically unstable [421] and allows to calculate the updated state |$\ket{\psi{(t_{0}+\Delta t)}}$ only for very small time steps $\Delta t$. However, this Taylor expansion indicates that the updated state |$\ket{\psi{(t_{0}+\Delta t)}}$ is a linear combination of vectors from the Krylov space $\mathcal{K}_{m}$ spanned by {|$\{\ket{\psi\left(t_{0}\right)},\hat{H}\ket{\psi\left(t_{0}\right)},\hat{H}^{2}\ket{\psi\left(t_{0}\right)},\dots,\hat{H}^{m-1}\ket{\psi\left(t_{0}\right)}\}$, which can be generated, to a given order $m$, by the Arnoldi algorithm [33]. Denoting by ${V_{m}}\in\mathbb{C}^{\mathcal{N}\times m}$ the matrix that maps the vectors of the canonical basis $\{\vec{e}_{i}\}$ in $\mathbb{C}^{m}$ to the naturally ordered vectors of the Krylov space, defines the step of the Krylov time evolution
@@
Both Chebyshev and Krylov time evolution algorithms utilize the sparse structure of the Hamiltonian matrix and allow for an efficient, numerically exact, calculation of the time-evolved state |$\ket{\psi(t)}$ under a many-body Hamiltonian $\hat{H}$. The range of system sizes available to those methods extends significantly beyond the capabilities of full ED. Computations performed on single nodes of present-day supercomputers allow to reach Hilbert space dimension $\approx 10^{9}$ [389, 673], corresponding to $L=32$ for the XXZ spin chain (7), and time scales of the order of $10^{3}$ tunneling times [590]. Implementation of the methods that employs internode communication allows to tackle time evolution for $\mathcal{N}\approx 10^{10}$, which corresponds to $L=36$ for the XXZ spin chain [89].
V.3 Tensor network approaches

### v3

@@ IV.4 Conclusion
In the following Section V, we briefly outline numerical approaches typically employed in studies of ergodicity and MBL. A reader interested in physics of disordered quantum many-body may skip Sec. V, and proceed directly to Sec. VI, in which we review numerical investigations of the ETH-MBL crossover in disordered many-body systems, providing concrete examples of the quantities that follow the general patterns described in the present Section. Subsequently, we focus on unitary dynamics of disordered many-body systems in Sec. VII and its relation to transport in Sec. VIII. In Table 1, we summarize the most important quantities employed in these considerations.
@@ V.1 Exact diagonalization
Exact diagonalization (ED) is a numerical method in which a certain number of eigenvalues $E_{m}$ and eigenvectors |$\ket{m}$ of Hamiltonian $\hat{H}$ of a many-body system is calculated. To that end, one needs to specify the matrix $\hat{H}$ corresponding to the Hamiltonian $\hat{H}$ in a selected many-body basis, as described, e.g., in an introductory fashion in [789, 329]. Once the Hamiltonian matrix $\hat{H}$ is calculated, one may employ a full exact diagonalization routine, using one of the available linear algebra packages, for instance, the LAPACK package [28], to compute all eigenvalues $E_{m}$ and eigenvectors |$\ket{m}$ of $\hat{H}$. Subsequently, the questions about thermalization (1) in the many-body system described by the Hamiltonian $\hat{H}$ can be answered. However, understanding whether the investigated system is in the MBL phase (6), or whether it follows the ETH, requires taking the large system size limit $L$. The computational resources required by the full ED quickly become prohibitive with increasing $L$ due to the exponential growth of the Hilbert space dimension $\mathcal{N}$ with the system size $L$, impeding the extrapolation of the results to $L\to\infty$. The typical matrix size accessible for full ED calculations in present-day computational studies is of the order of $\mathcal{N}\sim 10^{5}$. Hence, full ED calculations for the disordered XXZ spin chain (7) in the $\sum_{i}\hat{S}^{z}_{i}=0$ sector are restricted to $L\leq 18$. Note that investigations of MBL usually involve disorder average, which makes studies of larger system sizes infeasible.
@@
An observation that enables the exploration of system sizes beyond the reach of full ED is that the Hamiltonians of quantum many-body systems can often be written as sparse matrices for an appropriately chosen basis. The number of non-zero off-diagonal entries per row in such sparse matrices is much smaller than the matrix size $\mathcal{N}$. For example, the Hamiltonian of the disordered XXZ spin chain (7), written in the eigenbasis of $\hat{S}^{z}_{i}$ operators, contains no more than $L$ non-zero off-diagonal elements per row. By retaining the non-vanishing elements of the sparse Hamiltonian matrix $\hat{H}$ solely, it becomes possible to significantly reduce the computation time associated with matrix-vector multiplication in comparison to a dense matrix (which, by definition, has $O({\mathcal{N}})$ non-zero entries per row). This opens up the possibility of using the Lanczos algorithm [382], which employs repeated matrix-vector multiplications to compute eigenvectors corresponding to eigenvalues at the edges of the spectrum of $\hat{H}$. The Lanczos algorithm’s convergence slows considerably as the number of requested eigenvectors at the spectrum edges increases. The Lanczos algorithm needs to be paired with a so-called spectral transformation to make it practical to calculate the eigenvectors in the middle of the spectrum of $\hat{H}$ that are of interest in studies of MBL.
Shift-and-invert method relies on a transformation of the Hamiltonian matrix $\hat{H}$ via
$\hat{H}\rightarrow\hat{R}_{\sigma}(\hat{H})=(\sigma-\hat{H})^{-1},$ (21)
@@
where $\sigma$ is a chosen energy target. The spectral transformation (21) has two important properties: i) the eigenvectors |$\ket{m}$ of $\hat{H}$ are still eigenvectors of the transformed matrix $R_{\sigma}(\hat{H})$, ii) the eigenvalues are mapped according to $E_{m}\rightarrow R_{\sigma}(E_{m})$, which, by the construction of the function, ensures that the eigenvalues in the vicinity of the energy target $\sigma$ become the eigenvalues at the spectrum edges of the transformed matrix $R_{\sigma}(\hat{H})$. The two properties imply that the Lanczos algorithm for the transformed matrix $R_{\sigma}(\hat{H})$ converges to eigenvalues of $\hat{H}$ close to $\sigma$. Shift-and-invert method proved to be fruitful in studies of MBL [430], extending to $L\leq 26$ the range of system sizes for which calculation of eigenvectors in the middle of the spectrum of (7) is possible. The main difficulty of the shift-and-invert method is the presence of the inverse operator in (21). For interacting many-body systems, the Lanczos iteration with $R_{\sigma}(\hat{H})$ can be performed efficiently by a massively parallel $LU$ decomposition [26, 290, 44]. However, the sparsity pattern of the Hamiltonian matrix $\hat{H}$ is not preserved during the $LU$ decomposition, which leads to a huge memory overhead, and was identified as the main bottle-neck of the shift-and-invert method [523].
@@
Polynomial spectral transformations [587], which replace the complicated function (21) by a polynomial of $\hat{H}$, constitute an alternative to the shift-and-invert method. Calculation of eigenvectors in the middle of the spectrum of a large sparse matrix such as (7) can be efficiently carried out with an algorithm POLFED [646] that relies on the following polynomial spectral transformation,
$\hat{H}\rightarrow P^{K}_{\sigma}(\hat{H})=\frac{1}{D}\sum_{k=0}^{K}c^{\sigma}_{k}T_{k}(\hat{H}),$ (22)
@@
where $T_{k}(x)$ denotes $k$-th Chebyshev polynomial, the coefficients read $c^{\sigma}_{k}=\sqrt{4-3\delta_{0,k}}\cos(n\arccos{\sigma})$, and $D$ is a normalization factor. For sufficiently high and appropriately chosen order $K$, the polynomial $P^{K}_{\sigma}(E)$ has a sharp peak around the energy target $\sigma$ since the coefficients $c^{\sigma}_{n}$ are obtained from expanding a Dirac delta function centered at this $\sigma$ value. Hence, the Lanczos algorithm for $P^{K}_{\sigma}(\hat{H})$ converges to eigenvectors of $\hat{H}$ with energies around the chosen energy target. Notably, besides basic linear algebra operations, the Lanczos algorithm employs only the multiplication of vectors by $P^{K}_{\sigma}(\hat{H})$. The multiplication of vectors by $P^{K}_{\sigma}(\hat{H})$ reduces to repeated matrix-vector multiplications by the matrix $\hat{H}$ and linear algebra operations. Hence, the sparse structure of $\hat{H}$ is preserved by the POLFED algorithm, which significantly reduces the procedure’s memory cost. For instance, POLFED requires $\approx 100$ GB of RAM to calculate $2000$ eigenvectors in the middle of the spectrum of (7) for $L=24$, easily fitting into a single computing node of a present-day supercomputer. The shift-and-invert method requires $\approx 10^{4}$ GB of RAM for this task. At the same time, the total CPU time used by the two methods is of similar order; see [646] and [523] for further benchmarks. In passing, we note that combining a spectral transformation based on a second-order polynomial of $\hat{H}$ with appropriate preconditioning methods provides another exact diagonalization approach to MBL [56].
@@
Finally, the polynomial spectral transformations can also be used for efficient calculations of eigenvectors for large Floquet operator $\hat{U}_{F}$ [421]. In that case, analogously to the POLFED algorithm, one considers a high order polynomial of the Floquet operator,
$g_{K}(\hat{U}_{F})=\sum_{m=0}^{K}e^{-im\phi_{\mathrm{tg}}}\hat{U}_{F}^{m},$ (23)
@@
where $\phi_{\mathrm{th}}$ is the target eigenphase and the order of polynomial $K$ is fixed as $K=f\frac{\mathcal{N}}{N_{\mathrm{ev}}}$, where $N_{\mathrm{ev}}$ is the number of requested eigenvectors, the Hilbert space dimension is $\mathcal{N}=2^{L}$, and the factor $f=1.46$ was obtained from an optimization of the performance of the algorithm. The Arnoldi iteration [35] for the operator $g_{K}(\hat{U}_{F})$ yields $N_{\mathrm{ev}}$ eigenvectors |$\ket{\psi_{n}}$ corresponding to the dominant eigenvalues of $g_{K}(\hat{U}_{F})$ (which is no longer a unitary operator). The latter, by construction, correspond to eigenvectors of $\hat{U}_{F}$ with eigenphases closest to $\phi_{\mathrm{tg}}$. The convergence is reached after approximately $\alpha N_{\mathrm{ev}}$ steps of the iteration, where $\alpha\approx 2.1$. This algorithm is especially effective when the matrix-vector multiplication by the operator $\hat{U}_{F}$ can be efficiently implemented. This is the case for the Kicked Ising model, as described in Sec. X.4, or for various types of quantum circuits, one example was considered by [463]. For these cases, on present day computers, the algorithm enables study of eigenphases and eigenvectors of Floquet operators for $L=20$ or more (corresponding to matrix size $\mathcal{N}>10^{6}$), while the full exact diagonalization of Floquet operators is feasible up to $L\approx 15$ and $\mathcal{N}\approx 3\cdot 10^{4}$. In general, efficient matrix vector products can be formulated with a matrix product operator formulation of $\hat{U}_{F}$ [788].
V.2 Time evolution methods
@@
Exact diagonalization approaches allow us to study eigenvectors of $\hat{H}$, which encode, via (2), the limit of infinite times, $t\to\infty$. Moreover, the complete information about the eigenvalues and eigenvectors enables calculation of the time evolved state |$\ket{\psi(t)}$, and consequently, the time evolution of any observable $\hat{A}$ with (1). The memory complexity of the full ED is $O(\mathcal{N}^{2})$, while the space needed to store the time evolved state |$\ket{\psi(t)}$ scales linearly with $\mathcal{N}$. This observation, together with the sparsity of the Hamiltonian matrix $\hat{H}$, motivates the search for more effective time evolution algorithms.
@@
The efficient algorithm for time propagation of the Schrödinger equation was proposed in [688]. The algorithm is based on an expansion of the time evolution operator $\hat{U}(\Delta t)=e^{-iH\Delta t}$ over a time step $\Delta t$ into Chebyshev polynomials $T_{k}(x)$ of the Hamiltonian matrix
$\hat{U}(\Delta t)\approx\mathrm{e}^{-\mathrm{i}b\Delta t}\left(J_{0}(a\Delta t)+2\sum_{k=1}^{N}(-i)^{k}J_{k}(a\Delta t)T_{k}\left(\mathcal{H}\right)\right),$ (24)
@@
where $a=(E_{\rm max}-E_{\rm min})/2$, $b=(E_{\rm max}+E_{\rm min})/2$ and $E_{\rm min}$ ($E_{\rm max}$) is the energy of the ground state (the highest excited eigenstate) of $\hat{H}$. The Hamiltonian is rescaled as $\mathcal{H}=\frac{1}{a}(H-b)$, and $J_{k}(t)$ is the Bessel function of the order $k$. The order of expansion, $N\ll\mathcal{N}$, needed to assure convergence of (24) up to any desired precision, can be determined numerically for a given time step $\Delta t$, see, for instance, [653]. An application of $\hat{U}(\Delta t)$ allows to update the state of the system |$\ket{\psi{(t_{0}+\Delta t)}}=\hat{U}(\Delta t)\ket{\psi{(t_{0})}}$ and, upon repetition, results in the time-evolved state |$\ket{\psi(t)}$ with $t>t_{0}$. Since, for a fixed order $N$, $\hat{U}(\Delta t)$ is simply a polynomial of the matrix $\hat{H}$, the calculation of $\hat{U}(\Delta t)\ket{\psi{(t_{0})}}$ reduces to multiple multiplications of vectors by matrix $\hat{H}$ and linear algebra operators. Hence, the efficiency of the method relies crucially on the sparsity of the matrix $\hat{H}$. This Chebyshev time evolution algorithm allows to calculate the time-evolved state |$\ket{\psi(t)}$ couple of times faster than Suzuki-Trotter-based schemes [179], see also [212].
@@
A closely related method, called Krylov time evolution, was proposed in [482]. A direct truncation of the Taylor expansion of the time evolution operator $\hat{U}(\Delta t)\ket{\psi{(t_{0})}}=\sum_{k}\frac{(-i\Delta t)^{k}}{k!}H^{k}\ket{\psi{(t_{0})}}$ to order $m$ is numerically unstable [457] and allows to calculate the updated state |$\ket{\psi{(t_{0}+\Delta t)}}$ only for very small time steps $\Delta t$. However, this Taylor expansion indicates that the updated state |$\ket{\psi{(t_{0}+\Delta t)}}$ is a linear combination of vectors from the Krylov space $\mathcal{K}_{m}$ spanned by {|$\{\ket{\psi\left(t_{0}\right)},\hat{H}\ket{\psi\left(t_{0}\right)},\hat{H}^{2}\ket{\psi\left(t_{0}\right)},\dots,\hat{H}^{m-1}\ket{\psi\left(t_{0}\right)}\}$, which can be generated, to a given order $m$, by the Arnoldi algorithm [35]. Denoting by ${V_{m}}\in\mathbb{C}^{\mathcal{N}\times m}$ the matrix that maps the vectors of the canonical basis $\{\vec{e}_{i}\}$ in $\mathbb{C}^{m}$ to the naturally ordered vectors of the Krylov space, defines the step of the Krylov time evolution
@@
Both Chebyshev and Krylov time evolution algorithms utilize the sparse structure of the Hamiltonian matrix and allow for an efficient, numerically exact, calculation of the time-evolved state |$\ket{\psi(t)}$ under a many-body Hamiltonian $\hat{H}$. The range of system sizes available to those methods extends significantly beyond the capabilities of full ED. Computations performed on single nodes of present-day supercomputers allow to reach Hilbert space dimension $\approx 10^{9}$ [425, 727], corresponding to $L=32$ for the XXZ spin chain (7), and time scales of the order of $10^{3}$ tunneling times [639]. Implementation of the methods that employs internode communication allows to tackle time evolution for $\mathcal{N}\approx 10^{10}$, which corresponds to $L=36$ for the XXZ spin chain [92]. For specifically arranged many-body dynamics in which a finite number of subsystems interact only sporadically in time, but otherwise evolve independently, hybrid Schrödinger-Feynman techniques may offer an advantage [563] over the Chebyshev and Krylov time evolution algorithms.
V.3 Tensor network and other approaches